OpenStudy (anonymous):

another linear equation by substitution. x+3y=4 -x+2y=6

OpenStudy (anonymous):

dibs

OpenStudy (compassionate):

Dang flabbit @iambatman , first Eric and now you.

OpenStudy (anonymous):

Solve for x or y in one the equations, and then plug it into the second equation, and it will get rid of one of the variables so then you only have one variable to solve for. Once you solve for it, you can plug it back in any of the original equations and find the other :)

OpenStudy (anonymous):

i know its x=4y-3 but i always get confused with where to divide on both sides how do u know what sides to divide by

OpenStudy (anonymous):

come help me

OpenStudy (anonymous):

x=4-3y

OpenStudy (anonymous):

i mean x=4-3y

OpenStudy (anonymous):

sorry i got confused but yea explain that part to me

OpenStudy (anonymous):

Ok, so x = 4 - 3y -(4-3y)+2y=6

OpenStudy (anonymous):

Now solve for y

OpenStudy (anonymous):

You see we subbed it into the second equation, here I'll draw it out for you

OpenStudy (anonymous):

 can you continue?